OpenStudy (anonymous):

A ball is dropped from a height of 12m. Each time it strikes the ground it bounces back to a height of 75% the distance from which it fell. Determine the total distance traveled by the ball before it comes to rest.

OpenStudy (raden):

use this formula : total distance = h0 * (b+a)/(b-a) look at that 75/100 = 3/4 it mean b=4 and a=3 and given h0 = 12 m now, u have total distance = h0 * (b+a)/(b-a) = 12 * (4+3)/(4-3) = ...

OpenStudy (anonymous):

I think I have to use the inifinite sum series or something similar to that to solve this problem cause thats the chapter we're on :/

OpenStudy (raden):

hehe.. sorry, that's the shortcut formula :) but that will works

OpenStudy (anonymous):

ah okay. would you happen to know the other way?

OpenStudy (raden):

 total distance = 12 + 2(3/4*12 + 3/4*(3/4*12)+... so on) actually, in bracket is the geometry series